how many gallons of a 70% antifreeze solution must be mixed with 90 gallons of 25% antifreeze to get a mixture othat is 60% antifreeze

Answer by mananth(16946)

percent ---------------- quantity
Antifreeze I 25 ---------------- 90 gallons
Antifreeze II 70 ---------------- x gallons
Mixture 60 ---------------- 90 + x gallons

25 * 90 + 70 x =60(90+ x )
2250 + 70 x = 5400 + 60 x
70 x + -60 x = 5400 -2250
10 x = 3150
/ 10
x = 315
315 gallons of 70 % Antifreeze II